Alert Medical, LLC, consists of two doctors, Abrams and Lipscomb, who share in all income and losses according to a 2:3 income sharing ratio. Dr. Lin has been asked to join the LLC. Prior to admitting Lin, the assists of Alert medical were revalued to reflect their current market values. The revaluation resulted in medical equipment being increased by $28,000. Prior to the revaluation, the equity balances for Abrams and Lipscomb were $269,000 and 308,000.
media%2F768%2F768973a8-9184-450c-9f03-30
Answer #1

Part A

Account titles and explanation debit credit
Machinery 28000
Abrams member equity (28000*2/5) 11200
Lipscomb member equity (28000*3/5) 16800

Part B 2

Accounts title and explanation debit credit
Cash 175000
Abrams member equity (23750*2/5) 9500
Lipscomb member equity (23750*3/5) 14250
Lin member equity 151250

Lin member equity = (25%*((269000+11200)+(308000+16800)+175000) = 151250

Goodwill to Abram and Lipscomb = 175000-151250=23750